OpenStudy (anonymous):

f(x) = 3x and g(x) = 2x-6 find (f o g) (x)

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(2x-6)=3(2x-6)=6x-18\]

OpenStudy (anonymous):

Thank you~ <3